OpenStudy (anonymous):

2. What is the length of the hypotenuse of the right triangle shown? Round your answer to the nearest tenth if necessary. right triangle with sides 11 and 14, hypotenuse c (1 point) 10 17.8 625 5

OpenStudy (gorv):

\[\sqrt{11^2+14^2}\]

OpenStudy (gorv):

\[\sqrt{121+196}=\sqrt{317}\]

OpenStudy (gorv):

17.8 @ichliebediesetage

OpenStudy (anonymous):

Thank you. :)